### v1

@@ Abstract
The transaction pool plays a critical role in processing and disseminating transactions in cryptocurrency networks. However, increasing transaction loads strain the resources of full node deployments. We present Neonpool, an innovative transaction pool optimization using bloom filter variants, which reduces the memory footprint of the transaction pool to a fraction. Implemented in C++ and benchmarked using a unique Bitcoin and Ethereum dataset, our solution verifies and forwards transactions with over 99.99% accuracy and does not necessitate a hard fork. Neonpool is ideally suited for lightweight cryptocurrency clients and for resource-constrained devices such as browsers, systems-on-a-chip, mobile or IoT devices.
Index Terms:
Cryptocurrencies, full-node, IoT, lightweight, memory, mempool, optimization, txpool
@@ I Introduction
Increased transaction loads substantially increase the size of the pool, which strains the resources of nodes, and results in dropped transactions, processing delays, spikes in transaction fees, and even exposes the network to sophisticated attacks. Additionally, the transaction pool is also a vector for spam and dust attacks. In October 2015, a Bitcoin spam campaign grew the transaction pool to 1 GB (88,000 transactions), crashing 10% of nodes, mostly on Raspberry Pi [8].
@@
Neonpool utilizes two key insights: first, we observe that a majority of light clients levy a burden on the network, by piggybacking on existing full-node clients for their proper functioning. Second, we note that the two key functions of the transaction pool, inventory and forwarding, can be dissociated. This approach is similar to how lightweight clients (e.g. pruned nodes, SPV wallets) are commonly built, by prioritizing one function over another [3]. So we ask the question, is it possible to get the best of both worlds, i.e. design a client that optimizes local memory, and contributes to the health of the network, without imposing a significant burden on existing full-nodes?
@@
  • We describe Neonpool, an optimized transaction pool construction for cryptocurrencies that explores the use of standard bloom filters, decaying bloom filters, and bloom filters with key-value storage to replicate the transaction pool’s core function of transaction inventory. It reduces the local memory consumption of the transaction pool by up to two orders of magnitude (400 MB to 2 MB) while still processing unconfirmed transactions with over 99.99% fidelity.
@@
  • Neonpool-BTC and Neonpool-ETH are theoretically and empirically evaluated on multiple dimensions including error rates, memory utilization, computation time, and security on popular IoT devices.
To the best of our knowledge, Neonpool is the first optimization solution targeted specifically at the transaction pool and local memory. Our results demonstrate a dramatic reduction in memory consumption, up to 200x (400 MB to 2 MB for Bitcoin and Ethereum), with transaction processing accuracy over 99.99%. This solution enables resource-constrained systems like smartphones, systems-on-a-chip, mobile, and IoT devices to run a high-performing functional transaction pool. It does not require a hard fork and is orthogonal to other light clients. Neonpool may therefore be combined with them to aggregate their benefits. Neonpool can be extended to other cryptocurrencies.
Neonpool helps reduce the cost of running a full node for users. Running a network node contributes to the health of the network: it helps keep the network decentralized, as each node independently enforces consensus rules, and validates and verifies transactions. It also ensures privacy for users unlike SPV nodes and wallets that expose transaction history to external servers.
In the subsequent sections, we delve into the requisite background in §II, followed by the proposed scheme in §III. We analyze and discuss empirical results in §IV. We compare our scheme with prior work in §V. We identify potential future directions and conclude in §VI.
II Background
II-A Unconfirmed transaction pool
The unconfirmed transaction pool (alternately called the transaction pool) serves as a gateway for verifying and temporarily storing unconfirmed or pending transactions in a cryptocurrency node, while they await inclusion in a block.
Its primary functions include 1. Transaction verification: All incoming transactions are checked for adherence to the cryptocurrency’s protocol rules such as syntax, valid signatures, availability of funds, etc; 2. Transaction storage: A verified transaction is stored temporarily until it is included in a block. 3. Transaction propagation: The verified transactions are disseminated through the peer-to-peer network. A verified transaction is only forwarded to the peers the first time it is received by a node, to prevent loops in the network. Forwarding a transaction enables other nodes to independently verify and store them in their transaction pool.
Other functions include transaction fee estimation, transaction prioritization for inclusion in a block, and bootstrapping the transaction pool of a newly connected node.
@@ II-A Unconfirmed transaction pool
(a) Bitcoin [15]
(b) Ethereum [21]
Figure 1: Transaction count, raw size, and memory usage
However, map data structures have a notable drawback- they exhibit linear space complexity, and memory usage can be several times larger than the actual data size due to metadata, indexes, and pointers. We examine two major cryptocurrencies: Bitcoin, exemplifying a UTXO-based, and Ethereum, exemplifying an account-based cryptocurrency. Fig. 1(a)1(b), show the number of transactions, raw data size, and the resulting memory usage. Recently, Bitcoin and Ethereum’s transaction pool have frequently reached 400 MB in memory usage i.e. almost 3 times the size of raw data.
@@
II-B Bloom filter
A Bloom filter [30] is a probabilistic data structure used to test for set membership. It is particularly efficient in terms of memory usage but with a caveat: it may provide false positive results i.e. it can report that an element is in the set when it is not, but it will never produce a false negative.
Essentially it is a bit array of size m and k different hash functions. Initially, all bits are set to 0. To insert an element, it is hashed using each of the k hash functions, resulting in k different bits corresponding to positions in the bit array, which are set to 1. To check for membership, the element is again hashed by the k hash functions, and if any of the corresponding positions in the bit array are 0, the element is not in the set. If all the positions are 1, the element is likely in the set.
@@ II-B Bloom filter
In Figure 3 we insert three elements $x[1],x[2]$, $x[3]$ in the filter, which map to indices $\{2,9,15\}$, $\{6,10,15\}$$\{0,4,9\}$ respectively. For $x[4]$ with indexes $\{0,2,6\}$, the bloom filter returns a false positive due to earlier insertions setting those bits to 1. The probability of a false positive alternatively called the false positive rate (FPR) depends on the size of the bit array, the number of hash functions used, and the number of elements inserted. It is given by
@@
FPR highlights the characteristic trade-off between space and accuracy. Filter size $m$ can be provisioned as per set size $n$:
@@
The optimum value of a number of hash functions, $k$ is given by
@@
Decaying Bloom filters [37] randomly decay or ”age out” bits over time. This behavior is helpful when item relevance decreases over time, and it’s important to maintain an accurate representation of recent data while letting old data expire. Upon inserting an element into the Decaying Bloom filter, bits are reduced by a constant decay factor. A Decaying Bloom filter can be achieved by modifying a standard Bloom filter, whereupon every insertion $d$ random indices are decremented, mimicking expiry. For instance, the decay factor, $d=32$ results in 32 random indices of the filter being decremented, upon every insertion. However, a fraction of these randomly selected indices may be already zero.
III Proposed scheme
In this section, we provide a detailed explanation of our proposed scheme. Neonpool, does not necessitate the storage of complete transactions. Instead, it only stores transaction fingerprints, effectively disassociating the processes of transaction forwarding and inventory management.
Neonpool comprises two primary components. The first, bloomtxFilter, utilizes the transaction ID or hash, txHash to map valid ingress transactions. The second component, dstxFilter, ensures that duplicate or potential double-spend transactions are identified and discarded. The exact mechanism to identify potential double spending varies for UTXO-based Bitcoin and account-based Ethereum.
@@ III Proposed scheme
We present two variations of our proposed scheme, namely Neonpool-BTC and Neonpool-ETH. The term Neonpool is used if a certain aspect of the scheme applies to both Bitcoin and Ethereum. We use the term mempool to refer to the Bitcoin transaction pool, and txpool to refer to the Ethereum transaction pool, while transaction pool is a generic term that applies to both Bitcoin mempool and Ethereum txpool.
III-A Ingress
@@
Here, we describe the ingress process for the transaction pool and Neonpool, as depicted in Algo. 1 and 2 respectively.
@@
Figure 4: Neonpool-BTC Ingress (Where only (Y)es or (N)o is shown, the other implies the transaction is dropped.)
III-A1 Neonpool-BTC
@@
As shown in Fig. 4, in both Bitcoin Core and Neonpool-BTC the process begins with the arrival of a transaction announcement through an inv message. In Bitcoin, txHash is used to query the mempool to determine if the transaction already exists in the mempool. In Neonpool-BTC, the txHash is used to query the bloomtxFilter.
@@
Nodes may receive a transaction announcement multiple times but only accept it the first time they receive it. In both Bitcoin and Neonpool-BTC, if a transaction with the same txHash has already been received, it is discarded. If the transaction is determined to be new, the complete transaction is requested via a transaction message. When received, the complete transaction txBTC undergoes syntax, validity (valid transaction signatures, availability of sufficient funds etc), and semantics checks in both Bitcoin and Neonpool-BTC. Then the transaction is checked to detect double-spends.
@@
In both Bitcoin and Neonpool-BTC, each of the inputs, comprising the inputtxHash and index are scanned for double-spends. Transaction inputs are validated using the UTXO set. Transactions with invalid or spent inputs are discarded. It is also checked that none of the inputs exist in the mempool in Bitcoin. For Neonpool-BTC the dstxFilter is queried with the tuple <inputtxHash, index> to ascertain that the input does not already exist in the filter. If any of the inputs already exist in the mempool or dstxFilter, the transaction is dropped, as it constitutes a potential double spend. If two transactions with the same inputs are in circulation, the first seen by a node is regarded as safe, while the second is dropped. If any transaction input also (referred to as parent or ancestor) is missing, the transaction is added to the orphan pool. It will reside in the orphan pool until its ancestor is received, after which it will be re-processed. If the txBTC passes the verification process, it is added to the mempool in Bitcoin and bloomtxFilter in Neonpool-BTC. Finally, the transaction hash, txHash is relayed to the connected peers.
@@
Figure 5: Neonpool-ETH Ingress (Where only (Y)es or (N)o is shown, the other implies the transaction is dropped.)
III-A2 Neonpool-ETH
@@
As shown in 5, in both Ethereum and Neonpool-ETH a complete transaction txETH arrives via a transaction message. In Ethereum, the transaction hash (or ID) txHash is used to query the txpool to determine if the transaction already exists in the txpool. In Neonpool-ETH, the txHash is used to query the bloomtxFilter.
@@
Nodes may receive a transaction announcement multiple times but only accept it the first time they receive it. In both Ethereum and Neonpool-ETH, if a transaction with the same txHash has already been received, it is discarded. If the transaction is determined to be new, it undergoes syntax, validity (valid transaction signatures, funds availability etc), and semantics checks in both Ethereum and Neonpool-ETH.
@@
In both Ethereum and Neonpool-ETH, the transaction is then checked for potential double-spends by validating the transaction’s address, nonce and amount against the Trie. Transactions with an invalid or out-of-order nonce, or insufficient funds are rejected.

### v2

@@ Abstract
The increasing adoption of cryptocurrencies has significantly amplified the resource requirements for operating full nodes, creating substantial barriers to entry. Unlike miners, who are financially incentivized through block rewards and transaction fees, full nodes lack direct economic compensation for their critical role in maintaining the network. A key resource burden is the transaction pool, which is particularly memory-intensive as it temporarily stores unconfirmed transactions awaiting verification and propagation across the network. We present Neonpool, a novel optimization for transaction pool leveraging bloom filter variants to drastically reduce memory consumption by up to 200x (e.g., 400 MB to 2 MB) while maintaining over 99.99% transaction processing accuracy. Implemented in C++ and evaluated on unique Bitcoin and Ethereum datasets, Neonpool enables efficient operation on lightweight clients, such as smartphones, IoT devices, and systems-on-a-chip, without requiring a hard fork. By lowering the cost of node participation, Neonpool enhances decentralization and strengthens the overall security and robustness of cryptocurrency networks.
keywords:
@@ 1 Introduction
Increased transaction loads substantially increase pool size, which strains the resources of nodes, and results in dropped transactions, processing delays, spikes in transaction fees, and even exposes the network to sophisticated attacks. Additionally, the transaction pool is also a vector for spam and dust attacks. In October 2015, a Bitcoin spam campaign grew the transaction pool to 1 GB (88k transactions), crashing 10% of nodes, mostly on Raspberry Pi [7].
@@
Neonpool utilizes two key insights: first, we observe that a majority of light clients levy a burden on the network by piggybacking on existing full-node clients for their proper functioning. Second, we note that the two key functions of the transaction pool, inventory and forwarding, can be dissociated. This approach is similar to how lightweight clients (e.g., pruned nodes, SPV wallets) are commonly built by prioritizing one function over another [18]. So we ask the question, is it possible to get the best of both worlds, i.e., design a client that optimizes local memory and contributes to the network’s health without imposing a significant burden on existing full nodes?
@@
  • We describe Neonpool, an optimized transaction pool construction for cryptocurrencies that explores using standard bloom filters, decaying bloom filters, and bloom filters with key-value storage to replicate the transaction pool’s core function of transaction inventory. It reduces the transaction pool’s local memory consumption by up to two orders of magnitude (400 MB to 2 MB) while still processing unconfirmed transactions with over 99.99
@@
  • Neonpool-BTC and Neonpool-ETH are theoretically and empirically evaluated on multiple dimensions, including error rates, memory utilization, computation time, and security on popular IoT devices.
To the best of our knowledge, Neonpool is the first optimization solution targeted specifically at the transaction pool and local memory. Our results demonstrate a dramatic reduction in memory consumption, up to 200x (400 MB to 2 MB for Bitcoin and Ethereum), with transaction processing accuracy over 99.99%. This solution enables resource-constrained systems like smartphones, systems-on-a-chip, mobile, and IoT devices to run a high-performing functional transaction pool. It does not require a hard fork and is orthogonal to other light clients. Neonpool may be combined with them to aggregate their benefits. Neonpool can be extended to other cryptocurrencies.
Neonpool helps reduce the cost of running a full node for users. Running a network node contributes to the health of the network: it helps keep the network decentralized, as each node independently enforces consensus rules and validates and verifies transactions. It also ensures privacy for users, unlike SPV nodes and wallets, which expose transaction history to external servers.
In the subsequent sections, we delve into the requisite background in §2, followed by the proposed scheme in §3. We analyze and discuss empirical results in §4. We compare our scheme with prior work in §5. We identify potential future directions and conclude in §6.
2 Background
2.1 Unconfirmed transaction pool
The unconfirmed transaction pool (alternately called the transaction pool) serves as a gateway for verifying and temporarily storing unconfirmed or pending transactions in a cryptocurrency node while they await inclusion in a block.
Its primary functions include 1. Transaction verification: All incoming transactions are checked for adherence to the cryptocurrency’s protocol rules such as syntax, valid signatures, availability of funds, etc; 2. Transaction storage: A verified transaction is stored temporarily until it is included in a block. 3. Transaction propagation: The verified transactions are disseminated through the peer-to-peer network. A verified transaction is only forwarded to the peers the first time it is received by a node to prevent loops in the network. Forwarding a transaction enables other nodes to independently verify and store them in their transaction pool.
Other functions include estimating transaction fees, prioritizing transactions for inclusion in a block, and bootstrapping the transaction pool of a newly connected node.
@@ 2.1 Unconfirmed transaction pool
In Bitcoin, the mempool is allocated 300 MB by default [47]. In Ethereum, the default number of pending transactions in the txpool is 4096 [24]. Surplus transactions are evicted. Users can define a custom transaction pool acceptance policy. In a low-memory environment, the transaction pool size can be disabled entirely.
((a)) Bitcoin transaction count, raw size [17] and memory usage
((b)) Ethereum transaction count [29], raw size and memory usage
Figure 1: Transaction pool trends (since 2020)
Fig 1a provides the transaction count and raw transaction size of transactions in the Bitcoin mempool btccharts. Using this data, we calculate the memory usage of the Bitcoin mempool, which is approximately three times larger than the size of the raw transaction data [27] [26]. This disproportionate memory usage arises from the inherent limitations of map data structures, which require significant additional memory due to the storage of metadata, indexes, and pointers, resulting in total memory usage that far exceeds the size of the raw data. In early 2024, Bitcoin’s transaction pool often ranged from 150-300 MB in raw transaction size, with memory usage exceeding 400 MB.
Etherscan provides comprehensive data on transaction count for the Ethereum txpool [29], as shown in Fig 1b. As explained above, overheads are estimated to be three times the raw transaction size for Bitcoin. [^1] Simple ETH transfers are 30% of Ethereum transactions (210–250 bytes), while smart contract transactions make up 70% (500–1,000 bytes), resulting in an average size of  600 bytes. [28] This helps estimate memory usage. These metrics are visualized in Fig. 1b. For Ethereum, raw transaction size has frequently surpassed 150 MB, while memory consumption often crosses 400 MB.
@@
2.2 Bloom filter
A Bloom filter [38] is a probabilistic data structure used to test for set membership. It is particularly efficient with regards to memory usage but with a caveat: it may provide false positive results, i.e., it can report that an element is present in the set when it is not, but it will never produce a false negative.
Essentially it is a bit array of size m and k different hash functions. Initially, all bits are set to 0. To insert an element, it is hashed using each of the k hash functions, resulting in k different bits corresponding to positions in the bit array, which are set to 1. To check for membership, the element is again hashed by the k hash functions, and if any of the corresponding positions in the bit array are 0, the element is not in the set. If all the positions are 1, the element is likely in the set. In Fig. 3 we insert three elements $x[1],x[2]$, $x[3]$ in the filter, which map to indices $\{2,9,15\}$, $\{6,10,15\}$$\{0,4,9\}$ respectively. For $x[4]$ with indexes $\{0,2,6\}$, the bloom filter returns a false positive due to earlier insertions setting those bits to 1. The probability of a false positive, alternatively called the false positive rate (FPR), depends on the size of the bit array, the number of hash functions used, and the number of elements inserted. It is given by
@@ 2.2 Bloom filter
FPR highlights the trade-off between space and accuracy. Filter size $m$ can be provisioned as per set size $n$:
@@
The optimum value of a number of hash functions, $k$ is
@@
Decaying Bloom filters [46] randomly decay or "age out" bits over time. This behaviour is helpful when item relevance decreases over time, and it’s important to maintain an accurate representation of recent data while letting old data expire. Upon inserting an element into the Decaying Bloom filter, bits are reduced by a constant decay factor. A Decaying Bloom filter can be achieved by modifying a standard Bloom filter, whereupon every insertion $d$ random indices are decremented, mimicking expiry. For instance, the decay factor, $d=32$, results in 32 random indices of the filter being decremented upon every insertion. A fraction of these randomly selected indices may be already zero.
@@
Bloom filters in cryptocurrencies:Transaction Bloom filtering, introduced in BIP 37, enabled lightweight SPV wallets to efficiently retrieve transactions relevant to their addresses by sending a Bloom filter (based on its addresses, public keys, or other identifiers) to a full node, reducing bandwidth usage. However, BIP 37 was deprecated in Bitcoin Core 0.21.0 (2021) due to privacy concerns, as it exposed SPV wallet addresses to the full node. Graphene uses bloom filters to reduce network bandwidth in block reconciliation in Bitcoin [4]. In Ethereum, Bloom filters are used within block headers to summarize the logs generated by transactions in a block. This enables quick event log lookups without requiring full transaction execution. Bloom filters support block declaration fairness by enabling nodes to quickly verify transaction inclusion when competing blocks are declared, ensuring valid transactions are not overlooked. For a comprehensive summary of the applications of bloom filters in blockchain systems, we refer the reader to [5].
A Merkle tree is a binary tree structure used in cryptography to efficiently and securely verify the integrity of large datasets. They are invaluable for ensuring data integrity and inclusion in static, immutable datasets. Bitcoin and Ethereum leverage Merkle trees to organize transaction data within blocks. Once a block is mined and added to the chain, its data becomes immutable, making Merkle trees a suitable choice for this static context. However, their high update costs, verification overhead, and reliance on additional data (e.g., root and sibling hashes) make them unsuitable for the dynamic, high-churn environment of transaction pools.
In contrast, Bloom filters are particularly suited for optimizing the transaction pool because they enable efficient insertion and verification without requiring additional data, such as sibling or root hashes. Unlike Merkle trees, which scale logarithmically with the size of the dataset, the memory usage of a Bloom filter is independent of the dataset size. Instead, it is determined by its configuration, such as the size of the bit array and the number of hash functions. This makes Bloom filters especially beneficial for resource-constrained environments, including lightweight clients and IoT devices.
3 Proposed scheme
In this section, we provide a detailed explanation of our proposed scheme. By leveraging probabilistic data structures, Neonpool maintains the transaction pool’s core functionality while significantly improving resource efficiency. It does not necessitate the storage of complete transactions. Instead, it only stores transaction fingerprints, effectively disassociating the processes of transaction forwarding and inventory management.
Neonpool comprises two primary components. The first, bloomtxFilter, utilizes the transaction ID or hash, txHash, to map valid ingress transactions. The second component, dstxFilter, ensures that duplicate or potential double-spend transactions are identified and discarded. The exact mechanism to identify potential double-spending varies for UTXO-based Bitcoin and account-based Ethereum.
@@ 3 Proposed scheme
We present two variations of our proposed scheme, namely Neonpool-BTC and Neonpool-ETH. The term Neonpool is used if a certain aspect of the scheme applies to both Bitcoin and Ethereum. We use the term mempool to refer to the Bitcoin transaction pool and txpool to refer to the Ethereum transaction pool, while transaction pool is a generic term that applies to both Bitcoin mempool and Ethereum txpool.
3.1 Ingress
@@
Here, we describe the ingress process for the transaction pool and Neonpool, as depicted in Algo. 1 and 2, respectively.
((a)) Mempool Ingress
((b)) Neonpool-BTC Ingress
@@
Figure 4: Ingress: mempool vs Neonpool-BTC (only (Y)es or (N)o is shown, other implies transaction is dropped.)
3.1.1 Neonpool-BTC
@@
As shown in Fig. 4(a) and 4(b), in both Bitcoin Core and Neonpool-BTC, the process begins with the arrival of a transaction announcement through an inv message. In Bitcoin, txHash is used to query the mempool to determine if the transaction already exists in the mempool. In Neonpool-BTC, the txHash is used to query the bloomtxFilter.
@@
Nodes may receive a transaction announcement multiple times but only accept it the first time they receive it. This functionality is essential in cryptocurrency networks to minimize traffic overhead and prevent infinite loops by ensuring that transactions are broadcast only upon first receipt. In both Bitcoin and Neonpool-BTC, if a transaction with the same txHash has already been received, it is discarded. If the transaction is determined to be new, the complete transaction is requested via a transaction message. When received, the complete transaction txBTC undergoes syntax, validity (valid transaction signatures, availability of sufficient funds, etc.), and semantics checks in both Bitcoin and Neonpool-BTC, followed by checks to detect double-spends.
@@
In both Bitcoin and Neonpool-BTC, each of the inputs, comprising the inputtxHash and index are scanned for double-spends. Transaction inputs are validated using the UTXO set. Transactions with invalid or spent inputs are discarded. It is also checked that none of the inputs exists in the mempool in Bitcoin. For Neonpool-BTC the dstxFilter is queried with the tuple <inputtxHash, index> to ascertain that the input does not already exist in the filter. If any of the inputs already exist in the mempool or dstxFilter, the transaction is dropped, as it constitutes a potential double spend. If two transactions with the same inputs are in circulation, the first seen by a node is regarded as safe, while the second is dropped. If any transaction input also (referred to as parent or ancestor) is missing, the transaction is added to the orphan pool. It will reside in the orphan pool until its ancestor is received, after which it will be reprocessed. If the txBTC passes the verification process, it is added to the mempool in Bitcoin and bloomtxFilter in Neonpool-BTC. Finally, the transaction hash, txHash, is relayed to the connected peers.
((a)) Txpool Ingress
((b)) Neonpool-ETH Ingress
@@
Figure 5: Ingress: txpool vs Neonpool-ETH (only (Y)es or (N)o is shown, other implies transaction is dropped.)
3.1.2 Neonpool-ETH
@@
As shown in Fig. 5(a) and  5(b), in both Ethereum and Neonpool-ETH a complete transaction txETH arrives via a transaction message. In Ethereum, the transaction hash (or ID) txHash is used to query the txpool to determine if the transaction already exists in the txpool. In Neonpool-ETH, the txHash is used to query the bloomtxFilter.
@@
Nodes may receive a transaction announcement multiple times but only accept it the first time they receive it. In both Ethereum and Neonpool-ETH, if a transaction with the same txHash has already been received, it is discarded. If the transaction is determined to be new, it undergoes syntax, validity (valid transaction signatures, funds availability, etc.), and semantics checks in both Ethereum and Neonpool-ETH.
@@
In both Ethereum and Neonpool-ETH, the transaction is then checked for potential double-spends by validating the transaction’s address, nonce and amount against the Trie. Transactions with an invalid or out-of-order nonce or insufficient funds are rejected.